OpenStudy (destinykiara99):

Write the expression in standard form. five divided by quantity three minus fifteen i.

OpenStudy (destinykiara99):

\[5\div3-15i\]

OpenStudy (destinykiara99):

Answer Choices: five divided by seventy eight. - twenty five divided by seventy eight. i five divided by seventy eight. + twenty five divided by seventy eight. i - five divided by seventy eight. + twenty five divided by seventy eight. i - five divided by seventy eight. - twenty five divided by seventy eight. i

OpenStudy (destinykiara99):

@hartnn @hero @sleepyjess @imqwerty

OpenStudy (zzr0ck3r):

\(\dfrac{5}{(3-15i)}\dfrac{(3+15i)}{(3+15i)}=\dfrac{15+75i}{3*3-15^2i^2}\) knowing that \(i^2=-1\) can you finish this?